if angle A and angle B are complementry, and angle A=3x+4 and angle B=5x+6, find x.

Answer by checkley71(8403)

(3x+4)+(5x+6)=90
3x+5x+4+6=90
8x+10=90
8x=90-10
8x=80
x=80/8
x=10
(3*10+4)+(5*10+6)=90
30+4+50+6=90
34+56=90
90=90